OpenStudy (anonymous):

I need help with the Sign Chart of the Polynomial Equality. 2x^2-3x-5>0

OpenStudy (anonymous):

\[2x ^{2}-3x-5>0 \] \[2x ^{2}-3x-5=0\] (2x-5)(x+1)=0 2x-5= 0.........x=5/2 x+1=0...........x=-1 my problem is how you put this in the sign graph?

OpenStudy (anonymous):

you see what im saying?

OpenStudy (phi):

you are looking for (2x-5)(x+1) >0 this happens when the stuff on the left side is positive that means if both (2x-5) and (x+1) are positive or both (2x-5) and (x+1) are negative (two negatives give a positive)

OpenStudy (anonymous):

that doesn't explain how i get the positive in the left, the negative in the middle, and the postive on the right...

OpenStudy (phi):

First step is solve for 2x-5>0 and x+1>0 x > 5/2 and x>-1 both have to be true so x> 5/2 2nd step is solve for 2x-5<0 and x+1<0 x<5/2 and x<-1 both have to be true, so x<-1 those are the two conditions that solve your problem

OpenStudy (anonymous):

jesus this is confusing lool... can you somehow elaborate on what i do after those steps?

OpenStudy (anonymous):

because i already got em

OpenStudy (anonymous):

or is this the answer on how to properly place the signs on the chart

OpenStudy (anonymous):